# In situ measurement of slow solar wind emerging from a pseudostreamer: a conjunction study with Parker Solar Probe and Solar Orbiter

### v2

@@ 6 Conclusion
While the source regions of the slow solar wind still remain an open question, the novel in situ measurements that the new generation of spacecraft, PSP and Solar Orbiter, provide allow us to delve deeper into this question and study the nature of the solar wind. Combining the spacecraft instrumentation via conjunctions allows us to make detailed measurements of the solar wind evolution through comparing quantities known to decohere with radial distance, such as the Alfvénicity, with radially invariant quantities, such as the elemental composition and charge state ratios, which are fixed in the plasma at low coronal altitudes and therefore utilize the wider range of instrumentation available further from the Sun. In this paper, we outlined the modeling methods and measurements we use to study the coronal source origin of fast and slow solar wind and connect these methods to in situ plasma measurements. With PFSS and MHD modeling, we detect fast wind from deep in coronal holes and that the classical slow solar wind originates from a combination of HCS, pseudostreamer, and active region plasma as found by (Abbo et al., 2016; D’Amicis et al., 2021, and references therein), while the slow Alfvénic solar wind originates from a pseudostreamer stalk connected to a CH boundary. We then look at elemental abundances and particle measurements to study the FIP effect and trace coronal source region. In situ measurements from PSP and Solar Orbiter support our modeling results: fast solar wind has low FIP bias as measured by (Figure 4), very low ion charge state ratios implying a relatively cool coronal source, and higher alpha particle abundance ratios in comparison with the slow wind region, characteristic of wind streaming along open field lines from coronal holes. In contrast, slow solar wind shows a variance in density, abundance, and ion ratios implying a combination of source regions. We see both alpha-rich (CH origins) and alpha-poor (active region and/or streamer wind) slow wind and increased ionization states of carbon and oxygen in some slow wind regions, indicating wind from hotter, active region plasma. The ratio shows variance in the slow solar wind, oscillating between values characteristic of photospheric and coronal plasma.
In addition to providing evidence towards the source region of the plasma, the in situ measurements provide interesting insight into the plasma escaping from higher in the corona. There is a clear enhancement in charge state ratios and FIP bias as the spacecraft cross the HCS. This characteristic HCS signature seen in the SWA/HIS data shows that we are able to attach composition information to a PSP data set, extending its science potential and capabilities. During this encounter, we uncover a pseudostreamer over which we see an enhancement in the heavy ion ratios and (Figure 4) as measured by Solar Orbiter SWA/HIS, a jump in photospheric footpoints, increased proton densities, low Alfvénicity, and no crossing of the HCS. There is also a dip in the velocity differential as we cross this region and the lowest velocities measured during this time frame, supporting the idea of the slow solar wind partially originating from pseudostreamer regions that persists out into the heliosphere as both PSP and Solar Orbiter measured low velocities.
@@
We see very low alpha abundance in the FSW region and additional exploration of PSP fast wind periods should be done to understand if this is an effect of PSP’s radial position, this encounter specifically, or some other mechanism at work. The slow regions showed a large spread in source properties and coronal conditions for similar solar wind speed, from a priori very similar looking corona, and more work looking at the characteristics of the slow wind in comparison to its Alfvénicity and source region must be done to completely characterize the slow wind and its origination mechanisms. Additionally, further work exploring the impact of source region on wind temperature and the radial evolution of the parameters discussed in this paper (velocity, density, charge state ratios, FIP bias, helium abundance, Alfvénicity, etc.) should be done to fully characterize the solar wind propagating from various source regions. These types of studies require the existence and identification of future multi-spacecraft conjunctions to uncover the effects of radial propagation and have access to a variety of remote and in situ measurements to fully understand the processes at work creating and driving the solar wind.
@@ 7 Acknowledgements
This research used version 4.1.6 of the SunPy open source software package (Community et al., 2020), and made use of HelioPy, a community-developed Python package for space physics (Stansby et al., 2022). All code to replicate figures can be found at https://www.github.com/tamarervin/publications/e11_conjunction.
@@ 9 Appendix B: Magnetic Field Modeling
The top panel shows a radial cut at 13.03 of the radial velocity as modeled by MAS, along with the modeled HCS (in black). We see regions of faster wind emerging from between streamer arcs. In the middle panel, we show the radial cut of the density at 13.45 with density enhancements in the model corresponding to the HCS and streamer regions. In the bottom panel, we compare the model radial magnetic field at 13.45 with the measurements from PSP/FIELDS. We see in situ enhancements corresponding to the model enhancements and the PSP trajectory crossing pseudostreamer arcs around 165°, giving us another look at the model and a way to understand the coronal field at this time.

### v3

@@ 6 Conclusion
While the source regions of the slow solar wind still remain an open question, the novel in situ measurements that the new generation of spacecraft, PSP and Solar Orbiter, provide allow us to delve deeper into this question and study the nature of the solar wind. Combining the spacecraft instrumentation via conjunctions allows us to make detailed measurements of the solar wind evolution through comparing quantities known to decohere with radial distance, such as the Alfvénicity, with radially invariant quantities, such as the elemental composition and charge state ratios, which are fixed in the plasma at low coronal altitudes and therefore utilize the wider range of instrumentation available further from the Sun. In this paper, we outlined the modeling methods and measurements we use to study the coronal source origin of fast and slow Alfvénic solar wind and connect these methods to in situ plasma measurements. With PFSS and MHD modeling, we detect fast wind from deep in coronal holes while the slow Alfvénic solar wind originates from an over-expanded CH boundary.
We then look at elemental abundances and particle measurements to study the FIP effect and trace coronal source region. In situ measurements from PSP and Solar Orbiter support our modeling results: fast solar wind has low FIP bias as measured by (Figure 6), very low ion charge state ratios implying a relatively cool coronal source, and higher alpha particle abundance ratios in comparison with the slow wind region, characteristic of wind streaming along open field lines from coronal holes. The slow Alfvénic wind shows two plasma populations, one with FSW properties (low FIP, low charge state ratios) and another more similar to classical slow wind (higher charge state, higher FIP bias). In addition to providing evidence towards the source region of the plasma, the in situ measurements provide interesting insight into the plasma escaping from higher in the corona. There is a clear enhancement in charge state ratios and FIP bias as the spacecraft cross the HCS. This characteristic HCS signature seen in the SWA/HIS data shows that we are able to attach composition information to a PSP data set, extending its science potential and capabilities.
@@
We see very low alpha abundance in the FSW region and additional exploration of PSP fast wind periods should be done to understand if this is an effect of PSP’s radial position, this encounter specifically, or some other mechanism at work. The slow wind showed a large spread in source properties and coronal conditions for similar solar wind speed, from a priori very similar looking corona, and more work looking at the characteristics of the slow wind in comparison to its Alfvénicity and source region must be done to completely characterize the slow wind and its origination mechanisms. Additionally, further work exploring the impact of source region on wind temperature and the radial evolution of the parameters discussed in this paper (velocity, density, charge state ratios, FIP bias, helium abundance, Alfvénicity, etc.) should be done to fully characterize the solar wind propagating from various source regions. These types of studies require the existence and identification of future multi-spacecraft conjunctions to uncover the effects of radial propagation and have access to a variety of remote and in situ measurements to fully understand the processes at work creating and driving the solar wind.
@@ 7 Acknowledgements
The authors would like to thank the reviewer for their suggestions and feedback throughout the review process.
@@
This research used version 4.1.6 of the SunPy open source software package (Community et al., 2020), and made use of HelioPy, a community-developed Python package for space physics (Stansby et al., 2022). All code to replicate figures can be found at https://www.github.com/tamarervin/publications/e11_ch.
@@ 9 Appendix B: Magnetic Field Modeling
The top panel shows a radial cut at 13.03 of the radial velocity as modeled by MAS, along with the modeled HCS (in black). We see regions of faster wind emerging from between streamer arcs. In the middle panel, we show the radial cut of the density at 13.45 with density enhancements in the model corresponding to the HCS and streamer regions. In the bottom panel, we compare the model radial magnetic field at 13.45 with the measurements from PSP/FIELDS.